Find the inverse of the matrix using echelon reduction method .


x + 3y - z =0

x + y + z = 1

-x + 2y - z = 1


1
Expert's answer
2021-04-15T07:40:00-0400

Solution.

"A=\begin{pmatrix}
 1 & 3 &-1 \\
1&1&1\\
 -1&2&-1
\end{pmatrix}"

"A=\begin{pmatrix}
 1 & 3 &-1|1&0&1 \\
1&1 & 1 | 0&1&0\\
 -1&2&-1|0&0&1
\end{pmatrix}=
\begin{pmatrix}
 1 & 3 &-1|1&0&1 \\
0&-2&2|-1&1&0\\
 0&5&-2|1&0&1
\end{pmatrix}=\newline=
\begin{pmatrix}
 1 & 3 &-1|1&0&1 \\
0&1&-1|\frac{1}{2}&-\frac{1}{2}&0\\
 0&0&3| -\frac{3}{2}&\frac{5}{2}&1
\end{pmatrix}=
\begin{pmatrix}
 1 & 3 &-1|1&0&1 \\
0&1&-1|\frac{1}{2}&-\frac{1}{2}&0\\
 0&0&1| -\frac{1}{2}&\frac{5}{6}&\frac{1}{3}
\end{pmatrix}=\newline
=\begin{pmatrix}
 1 & 3 &0|\frac{1}{2}&\frac{5}{6}&\frac{1}{3} \\
0&1&0|0&\frac{1}{3}&\frac{1}{3}\\
 0&0&1| -\frac{1}{2}&\frac{5}{6}&\frac{1}{3}
\end{pmatrix}=
=\begin{pmatrix}
 1 & 0 &0|\frac{1}{2}&-\frac{1}{6}&-\frac{2}{3} \\
0&1&0|0&\frac{1}{3}&\frac{1}{3}\\
 0&0&1| -\frac{1}{2}&\frac{5}{6}&\frac{1}{3}
\end{pmatrix}."

So,

"A^{-1}=\begin{pmatrix}
 \frac{1}{2}&-\frac{1}{6}&-\frac{2}{3} \\
0&\frac{1}{3}&\frac{1}{3}\\
 -\frac{1}{2}&\frac{5}{6}&\frac{1}{3}
\end{pmatrix}."